OpenStudy (anonymous):

An equilateral triangle with side( 2 radical 3 ) of is inscribed in a circle. What is the area of one of the sectors formed by the radii to the vertices of the triangle? a) 1.33pi sq. in. b) 2pi sq. in. c) 2.09pi sq. in.

OpenStudy (anonymous):

it is not clear to me what sector they want is there a picture?

OpenStudy (anonymous):

lol yeah i got that !

OpenStudy (anonymous):

this is rather involved are you supposed to know trig or not?

OpenStudy (anonymous):



OpenStudy (anonymous):

That's supposed to be 2 radical 3

OpenStudy (anonymous):

yeah i got it

OpenStudy (anonymous):

and no they don't expect us to know trig at all

OpenStudy (anonymous):

here is a brief explanation the first part takes some work but i will just say it

OpenStudy (anonymous):

I've only been through algebra 1

OpenStudy (anonymous):

you need the radius of the circle that is 2

OpenStudy (anonymous):

so the area of the circle is \(4\pi\)and your sector is one third of that therefore \[\frac{4\pi}{3}\]

OpenStudy (anonymous):

how you get the radius of 2 is the real question i am trying to think of a non - trig explanation

OpenStudy (anonymous):

Oh Boy

OpenStudy (anonymous):

aww man it's Thundering outside ad raining. Ughhh

OpenStudy (anonymous):

here too

OpenStudy (anonymous):

you in florida right?

OpenStudy (anonymous):

You live in Florida too/

OpenStudy (anonymous):

no philly

OpenStudy (anonymous):

Wow! Lol!

OpenStudy (anonymous):

That kinda sucks, but still wow

OpenStudy (anonymous):

I live in Branford

OpenStudy (anonymous):

about an hour and a half away from Gainesville

OpenStudy (anonymous):

that is just a link to odysseyware

OpenStudy (anonymous):

Yeah, but it's the lesson they gave me

OpenStudy (anonymous):

i don't have an account, so i am just getting the log in page

OpenStudy (anonymous):

Does it show on your end

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

no problem if you want help with a question and it has a picture, post a screenshot

OpenStudy (anonymous):

I appreciate the answer anyway. I didn't see any more like that

OpenStudy (anonymous):

I'm on a computer though. How would I be able to do that/

OpenStudy (anonymous):

Wait. How about I just grant you access to my google docs. i have the lesson stuff on there all laid out

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

hit "print screen" on the keyboard save the picture then "attach file" below

OpenStudy (anonymous):

Oh. Thanks

OpenStudy (anonymous):

Did you still want to do it on email or keep it on here so you can get medals?

OpenStudy (anonymous):

i don't need the medal but it is easier to write math here

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Well that didn't turn out at all

OpenStudy (anonymous):

that didn't work

OpenStudy (anonymous):

Oh boy

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

not sure why

OpenStudy (anonymous):

Eh. Me neither

OpenStudy (anonymous):

I do admit it was kinda funny Lol!

OpenStudy (anonymous):

It was sooo TINY

OpenStudy (anonymous):

did you hit "print screen?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Let me try again and see what happens

OpenStudy (anonymous):

Still didn't work jeez

OpenStudy (anonymous):

\[\frac{70^\circ}{360^\circ}=\frac{7}{\pi r^2}\]

OpenStudy (anonymous):

I'm here now

OpenStudy (anonymous):

Sorry